# Robust fuzzy clustering with cellwise outliers

### v1

@@ 2.1 An EM-type algorithm for cellFCLUST
At the end of Step 4, the objective function in (4) is computed. If the focus is on clusters with $\sum_{i=1}^{n}u_{i1}^{m}=\ldots=\sum_{i=1}^{n}u_{iK}^{m}$, the objective function to be considered is (1), which approximately corresponds to set $p_{k}=1/K$ in Step 4 throughout all iterations (this is an exact correspondence for $m=1$). Steps 1-4 are repeated until a sensible increase in the objective function occurs, by setting an arbitrary small tolerance value $\epsilon$ (e.g., $10^{-6}$ in our experiments), or until a maximum number of iterations ($500$ in our cases) is reached. To increase the chances of finding the global constrained maximum of the objective function, the algorithm should be run several times with different initializations, retaining the best solution.
3 Effects of the tuning parameters: numerical examples
The methodology introduced in the previous section depends on several tuning parameters: the number of clusters ($K$), the contamination level ($\alpha$), the constant for the eigenvalue-ratio constraint ($c$), the fuzzifier ($m$), and the scale factor ($S$). The latter affects the solution of cellFCLUST by controlling the proportion of hard assignments. Given the property of high contrast, we treat the scale factor as a key tuning parameter that allows us to manage the mentioned level of hard assignment depending on the real case study. We illustrate herein the role of all tuning parameters in cellFCLUST through numerical examples on artificial data.
Figure 1: Artificial data with $5\%$ of contamination per variable. Stars denote observations that have outlying values in at least one of the first two variables
First, we generate $200$ units grouped into two clusters ($70\%$ in the first cluster, $30\%$ in the second one), measured on $5$ variables. For each variable, $5\%$ of the cells are randomly selected and contaminated. Specifically, for the first two variables – on which we will focus on for illustrative reasons – we consider the following contamination scheme: $4$ observations (units 9, 10, 15, 60) are unreliable in both variables with respect to the clustering structure; $12$ observations (units 28, 61, 71, 77, 103, 127, 144, 154, 156, 176, 190, 196) has only one cell that is contaminated (Figure LABEL:fig:_example_orig). For the other three variables, $5\%$ of the cells are replaced with uniform random values in [50,50][-50,50]. It is worth highlighting that the presence of these three additional variables is crucial for retaining reliable information used during the imputation step (Step 3), especially for observations contaminated in both of the first two variables. cellFCLUST applied on these data with $K=2$, $\alpha=0.05$, and $c=80$ correctly identifies the outlying values and the clustering structure, as shown in Figure LABEL:fig:_example_cellFClUST. The constant for the eigenvalue-ratio constraint is fixed considering that $\boldsymbol{\mathbf{S}}_{1}=[s_{ij}=0.9^{\rvert i-j\lvert}:i,j=1,\ldots,J]$ and $\boldsymbol{\mathbf{S}}_{2}=[s_{ij}=-0.8^{\rvert i-j\lvert}:i,j=1,\ldots,J]$, where $J=5$.
Figure 2: Artificial data: objective function curves
@@ 3 Effects of the tuning parameters: numerical examples
Number of clusters $K$ and contamination level $\alpha$: the number of clusters in a data set cannot be chosen independently of the level of contamination. In this framework, we evaluate the behavior of the curves representing the objective function in (4) at convergence by running cellFCLUST with different values of $K$ and $\alpha$, when $c=80$ and $m=1.5$. Assuming the data set is contaminated, we vary $\alpha$ over the set $\{0.01,0.05,0.10,0.20\}$. In this analysis, the cellFCLUST objective function for $\alpha=0$ (i.e., when no contamination is considered) behaves similarly to the case $\alpha=0.01$ across the values of $K$ in $\{1,\ldots,4\}$, and performs worse than higher levels of contamination. Therefore, we omit it in this specific example, although the user may wish to include this level when no prior information about the presence of outlying values in the data is available. As shown in Figure 2, the objective functions differ significantly when $\alpha=0.01$, since underestimating the contamination level requires additional clusters beyond those originally generated. On the other hand, when $\alpha=0.05$, which is the theoretical contamination level in the data set, the objective functions for $K=2,3,4$ are close to each other, indicating that two is a suitable choice for the number of clusters. Indeed, $K=2$ captures the underlying structure of the data well, and increasing this number does not lead to a meaningful improvement in the objective function.
@@
Figure 3: Artificial data: difference between the knee point of $\Delta_{ij}$ and the $\alpha$ value. The curve shows the median differences across variables as $\alpha$ varies, with $K=2$. The shaded gray area represents the variability
@@
Figure 4: Artificial data: $\Delta$ plots for the first two variables, where units are sorted according to their $\Delta$ values. Vertical, red line corresponds to the value of $\alpha$ used for the model implementation when $K=2$
@@
We further investigate how the cellFCLUST results change as the contamination level $\alpha$ varies, to support the previous finding by analyzing the behavior of $\Delta_{ij}$. These are obtained as defined in (6) using the parameters estimated at convergence for the selected $K$. For each variable, we compute the maximum distance between each $\Delta_{ij}$ and the straight line connecting the endpoints $\Delta_{(1)j}$ and $\Delta_{(n)j}$, while varying $\alpha$. This procedure corresponds to identifying the knee point of the $\Delta$ curve. Figure 3 shows the median of the difference between the identified knee point and the corresponding $\alpha$ value across the five variables. The difference approaches zero when $\alpha=0.05$, with the smallest variability (gray area, calculated using the median absolute deviation [25]), confirming the selection of this value as the appropriate contamination level for these data. Furthermore, we can verify the behavior of $\Delta_{ij}$ for each variable, recalling that they correspond to the difference between the individual contribution to the objective function when a cell is considered reliable or unreliable. In Figure 4, we plot points $\{\big{(}(i)/n,\Delta_{(i)j}\big{)}\}_{i=1}^{n}$ for the first two variables, where observations are ordered according to their $\Delta$ values. When $\alpha=0.01$, some observations with a small value of $\Delta$ are not considered as contaminated, resulting in a lower contamination level than the theoretical one (Figure LABEL:fig:_example_Delta_1). On the hand, too many cells than needed are flagged as unreliable when $\alpha=0.10$, as shown in Figure LABEL:fig:_example_Delta_10, where the values of $\Delta$ level off before the considered $\alpha$. The appropriate choice turns out to be $0.05$ (Figure LABEL:fig:_example_Delta_5), which corresponds to the proportion of contaminated cells in the artificial data. Indeed, at this level of contamination, $5\%$ of the $\Delta$ values per variable are significantly lower than the majority of the others. The same occurs for the last three variables, as reported in the Supplementary Material. In general, we recommend initially adopting a conservative choice of $\alpha$, then gradually decreasing it while monitoring the behavior of the objective function curves. The additional tools provided for supporting the choice of $\alpha$ are particularly useful when the objective function curves are smoother.
Figure 5: Artificial data: clustering structure with $K=2$ for two levels of $c$. Stars denote observations that have outlying values in at least one of the $5$ variables
Constant for the eigenvalue-ratio constraint $c$: the constant for the eigen-value-ratio constraint in (3) allows for different cluster shapes. When $c=1$, the clusters become spherical, and cellFCLUST with $\alpha=0$ produces similar results to those of fuzzy $c$-means. However, the cellwise approach to outlier detection can be affected by the choice of $c$, not only in terms of cluster shape but also in the identification of the outlying cells. In Figure 19, we illustrate this effect by highlighting the observations that have at least one unreliable cell among the five variables – unlike in Figure 1, where the stars represent observations with outlying values considering the first two variables. For instance, in Figure LABEL:fig:_example_c1, where $c=1$, units $21$ and $44$ are flagged as observations with outlying values, whereas they turn out to be fully reliable when $c=80$ (Figure LABEL:fig:_example_c80); the opposite occurs for unit $22$. By enforcing a spherical shape for the clusters, the constant $c$ constrains the estimation of the covariance matrices, consequently affecting the values of $\Delta$. This leads to a different configuration of the matrix $\mathbf{W}$, and consequently, to a different set of observations that have unreliable cells. It is worth noting that the data are five-dimensional, so the cluster configuration depends not only on the first two variables but also on the other three, even though they cannot be directly visualized in the full feature space. In this example, the clusters are well-separated and a wide range of $c$ gives similar results in terms of clustering (i.e., stable solutions). However, misclassifications may occur when clusters are not well-separated and $c$ is constrained to its smallest value (see the Supplementary Material).
The example considered so far has been used to illustrate the effect of the tuning parameters on the identification of the outlying cells and the cluster recovery. To generate inliers – such as observations $9$ and $10$ in Figure LABEL:fig:_example_orig, which have outlying cells for the first two variables and lie between the two clusters – the clustering configuration is well-separated. Therefore, the fuzzifier value $m=1.5$ has not a huge effect on the fuzzification, as expected. Considering a similar setup as before, we move the vectors $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ (cluster centers) closer together and restrict the number of variables to three, which is the maximum dimensionality that allows for complete visualization. Additionally, we contaminate only $5\%$ of the cells in the first variable, and we set $K=2$, $c=50$ and $\alpha=0.05$. Following the previously mentioned suggestion, a conservative choice for $\alpha$ is recommended and is not particularly detrimental even when the true contamination level is lower, as is the case for the second and third variables.
Figure 6: Artificial data: effect of the fuzzifier parameter $m$ on the cluster membership values. Stars denote observations that have outlying values in the first variable
[…]
Figure 7: Artificial data: effect of the fuzzifier parameter $m$ and the scale factor $S$ on the cluster membership values. Stars denote observations that have outlying values in the first variable
Interaction between the fuzzifier parameter $m$ and the scale factor $S$: similarly to its casewise counterpart, cellFCLUST shows different behavior depending on the scaling of the data, that is when each variable is scaled by a constant factor $S$. This occurs when $m>1$, whereas the results of cellFCLUST are scale-independent when $m=1$. Therefore, the scale factor and the fuzzifier interplay, resulting in a change in the clustering structure estimated by cellFCLUST when both $S$ and $m$ vary. Figure 7 provide evidence of this effect. Specifically, as the scale factor decreases and the fuzzifier increases, the assignments tend to become less crisp and more fuzzy, with memberships approaching equality across clusters. However, complete fuzzification is not desirable from an interpretation point of view. In this example, the outlying values in the first variable are correctly recognized in both Figures 6 and 7.
Figure 8: Artificial data: clustering results with $K=3$, $\alpha=0.05$, $S=2$, $m=1.9$, and unequal weights (left panel) or equal weights (right panel)
Restriction on the weights $p_{k}$: one of the advantages of cellFCLUST is the possibility to accommodate different cluster sizes, specifically through distinct cluster weights. To clarify the effect of the restriction $p_{k}=1/K$ for $k=1,\ldots,K$, ideally hypothesizing that this corresponds to $\sum_{i=1}^{n}u_{i1}^{m}=\ldots=\sum_{i=1}^{n}u_{iK}^{m}$, we analyze the previous example by changing the number of clusters to three in running cellFCLUST. We remind that $70\%$ of the observations are generated from the first (red) cluster, and $30\%$ from the second (green) one. In the left panel of Figure 8, we can see that a few observations are classified in the third cluster, and their assignments are extremely fuzzified to the first cluster (points in the figure inclined toward violet). Indeed, the weight estimate for the third cluster is $0.03$, and therefore the clustering structure appears to be very similar to the original one, despite the higher $K$. On the other hand, when the weights are constrained to be equal across clusters, the original first cluster is split into two of equal size, as shown in the right panel of Figure 8.
[…]

### v2

@@ 2.1 An EM-type algorithm for cellFCLUST
At the end of Step 4, the objective function in (4) is computed. If the focus is on clusters with $\sum_{i=1}^{n}u_{i1}^{m}=\ldots=\sum_{i=1}^{n}u_{iK}^{m}$, the objective function to be considered is (1), which approximately corresponds to set $p_{k}=1/K$ in Step 4 throughout all iterations (this is an exact correspondence for $m=1$). Steps 1-4 are repeated until a sensible increase in the objective function occurs, by setting an arbitrary small tolerance value $\epsilon$ (e.g., $10^{-6}$ in our experiments), or until a maximum number of iterations ($500$ in the examples shown in this work) is reached. To increase the chances of finding the global constrained maximum of the objective function, the algorithm should be run several times with different initializations, retaining the best solution.
3 Simulation study
We conduct a simulation study to evaluate the performance of cellFCLUST in cluster recovery, parameter estimation, and outlier detection compared to some alternative methodologies. The first competitor is F-TCLUST, which represents the casewise counterpart of our proposal and it is run via the R package tclust ([29], R version 3.3). Two other classes of methods encompass those which are not specifically tailored for outliers: fuzzy $k$-means (FKM, R function FKM) and Gustafson, Kessel and Babuska-like fuzzy $k$-means (FKMGKB, R function FKM.gkb), which relaxes the spherical assumption for the clusters [30]. Their robust versions based on the identification of a noise cluster are implemented via the R functions FKM.noise and FKM.gkb.noise, respectively. All these methods are included in the R package fclust [31]. Additionally, we also consider the Unsupervised Fuzzy Trimmed C Prototypes (UFTCP, [17]), where fuzzy $c$-means is extended via a trimming approach.
Two scenarios with varying numbers of observations and clusters are considered, while the number of variables is fixed to $J=10$. In Scenario 1, $n=250$, $G=2$: $30\%$ of observations are generated from the first cluster, with $\boldsymbol{m}_{1}=\mathbf{0}$ and $\boldsymbol{\mathbf{S}}_{1}=[s_{jl}=(0.6^{\lvert j-l\rvert})/16,j,l=1,\ldots,J]$; and $70\%$ from the second cluster, with $\boldsymbol{m}_{2}=[m_{j}=(-1)^{j}\times 0.5,j=1,\ldots,J]$ and $\boldsymbol{\mathbf{S}}_{2}=[s_{jl}=((-0.6)^{\lvert j-l\rvert})/16,j,l=1,\ldots,J]$. In Scenario 2, $n=500$, $G=4$, and the sizes of the clusters are more balanced: the first two clusters, each containing $20\%$ of the observations, have the same mean vectors and covariance matrices as in Scenario 1; the other two clusters, each composed of $30\%$ of the observations, have mean vectors $\boldsymbol{m}_{3}=[m_{j}=(j\mod 3)-1,j=1,\ldots,J]$ and $\boldsymbol{m}_{4}=[m_{4}=\big((j+1\mod 4)-1\big)/2,j=1,\ldots,J]$, and covariance matrices $\boldsymbol{\mathbf{S}}_{3}=[s_{jl}=(0.7^{\lvert j-l\rvert})/16,j,l=1,\ldots,J]$ and $\boldsymbol{\mathbf{S}}_{4}=[s_{jl}=\big((-0.7)^{\lvert j-l\rvert}\big)/16,j,l=1,\ldots,J]$. The maximum overlap between pairs of clusters – where the overlap is defined in [32] as the sum of the two misclassification probabilities – ranges from $0.01$ to $0.05$ in both scenarios, indicating moderate cluster separation, which represents a suitable configuration for fuzzy clustering methodologies. For each scenario, $100$ random samples are obtained and contaminated with $0\%$ (baseline), $1\%,5\%$, and $10\%$ of outlying cells per variable (i.e., in each column of the data matrix) randomly drawn from a uniform distribution on the interval [30,30][-30,30]. The contamination is also performed in such a way that no outlying observation lies within the $99$th percentile ellipsoid of any component, comparing their Mahalanobis distance, computed using the parameters employed for data generation, with the $99$th percentile of a chi-squared distribution with $J$ degrees of freedom.
The fuzzifier parameter $m$ is set to $2$ for cellFCLUST and F-TCLUST, to $1.35$ for FKM, FKM.noise and UFTCP, and to $1.5$ for FKM.gkb and FKM.gkb.noise. To select these values, we compare the proportion of Weak Assignments (WA) – defined as the proportion of observations whose highest membership is below $0.90$ – resulting from the estimation of each methodology under $\alpha=0$ (baseline) with the theoretical ones. The latter are computed from the membership matrices in (7) using the generating parameters and $\alpha=0$, yielding WA proportions of $0.05$ and $0.07$ in the two scenarios, respectively. It should be noted that we focus on $u_{ik}$ for the observation $i$ belonging to cluster $k$ rather than the fuzziness itself, i.e., $u_{ik^{\prime}}$ with $k^{\prime}\neq k$, since the latter depends on the number of clusters considered. Additionally, analyzing the highest membership value across clusters provides information on the “weight” an observation has in the cluster it is assigned to: if it is a representative observation of that cluster, its membership will be high – potentially approaching one – regardless of the number of clusters. In our experience, when an observation is weakly assigned to a cluster, the remaining membership is typically shared among only a few clusters, that is, it is not evenly split across all K1K-1 clusters. For the methods that require eigenvalue-ratio constraints given by $c$, i.e., cellFCLUST and F-TCLUST, this is set according to the ratio of the eigenvalues computed from the true (i.e., data-generating) covariance structure. For FKM.gkb and FKM.gkb.noise, we specify precise constraints: the volume parameters $\rho_{g}$ as obtained from the aforementioned covariance matrices, and $\gamma=0.1$ to ensure numerical stability. These two parameters are therefore not set to their default values, which were fixed only to prevent numerical singularities. Instead, we apply fine-tuning to choose more appropriate values that improve the performance of the clustering methods. The contamination level $\alpha$ is fixed to the true generating contamination for cellFCLUST and to $0.25$ for F-TCLUST, as previously employed in [22]. However, due to the nature of cellwise contamination, with $J=10$ and cellwise outliers representing $1\%$, $5\%$, and $10\%$ of the cells, the percentage of cases affected by at least one outlying cell can reach $10\%$, $40\%$, and $60\%$ of the total cases, respectively. For noise clustering methodologies, the number of outliers is automatically selected within the algorithm. Similarly, UFTCP chooses the appropriate outlier proportion from a set of possible levels ranging from 0 to 0.5, in increments of 0.05, ultimately retaining the value that maximizes a density criterion [8] as the validity measure.
Figure 2: Results of the simulation study: Misclassification Rate (MR) and Root Mean Squared Error (RMSE) of the membership matrix averaged over $100$ samples per scenario, percentage of contamination, and model. For F-TCLUST, FKM.noise, FKM.gkb.noise, and UFTCP the indices are computed only on the observations not flagged as outliers.
As can be seen in Figure 2, FKM, FKM.gkb, and UFTCP show the highest Misclassification Rate (MR) when the cellwise contamination is set to $1\%$, whereas at $5\%$ and $10\%$, only cellFCLUST and FKM.gkb.noise maintain good clustering performance. Specifically, at $10\%$, all models except cellFCLUST and FKM.gkb.noise exhibit extremely poor clustering recovery. This is mainly due to the large number of observations containing outlying cells and to the ability of these two models to accommodate non-spherical clusters. Moreover, unlike F-TCLUST, where $\alpha$ is fixed in advance, FKM.gkb.noise automatically selects the proportion of outlying observations, allowing it to stabilize at lower MR values than the proposal’s natural casewise counterpart. A similar behavior can be observed for the Root Mean Squared Error (RMSE) of the membership matrices. Besides, cellFCLUST always outperforms FKM.gkb.noise and this can be explained by the higher efficiency of the former when dealing with cellwise outliers, without discarding or downweighting the entire contaminated observations during the estimation procedure.
Figure 3: Results of the simulation study: Root Mean Squared Error (RMSE) of the component mean vectors and Kullback-Leibler (KL) discrepancy for the component covariance matrices averaged over $100$ samples per scenario, percentage of contamination, and model. The values are represented via log-transformation, while the y-axis ticks are labeled using the original scale.
To assess the parameters’ estimates, we solve the label switching problem by minimizing RMSE between the estimated means and the theoretical ones over the different possible permutations. As shown in Figure 3, RMSE of the component mean vectors follows a similar behavior to that of the membership matrices, already deteriorating at $1\%$ of contamination for the non-robust models. Additionally, it significantly degrades at $5\%$ and $10\%$ for F-TCLUST and FKM.noise, whereas the FKM.gkb.noise estimates remain acceptable but suboptimal with respect to cellFCLUST. For the Kullback-Leibler (KL) discrepancy of the component covariance matrices, cellFCLUST outperforms the competitors from the first contamination level. On the other hand, F-TCLUST provides good estimates only when the number of contaminated observations is less than those trimmed, handling at most $1\%$ of cellwise contamination when $J=10$. Unlike the other indices, the difference between cellFCLUST and FKM.gkb.noise becomes evident in the estimation of the covariance matrices. Indeed, for FKM.gkb-type models – especially the noise version – increasing the smallest eigenvalue of the covariance matrix $\mathbf{S}_{k}$ to ensure its determinant equals $\rho_{g}$ does not yield good results compared to the efficient procedure proposed in [28], which provides a closed-form solution for the truncated eigenvalues satisfying constraint (3). This is particularly pronounced in Figure 3 at $1\%$ contamination, where the two models implementing the aforementioned procedure, i.e., cellFCLUST and F-TCLUST, produce better estimates of the covariance matrices – F-TCLUST is implemented with $\alpha=0.25$, and the maximum percentage of observations with at least one contaminated cells is $10\%$ when the true $\alpha=0.01$. At higher contamination levels, the KL discrepancy of FKM.gkb.noise is lower than that of F-TCLUST because of the larger proportion of outliers not flagged by the latter. Regarding cellFCLUST, both the effect of the constraint and the cellwise outlier detection allow this methodology to achieve better KL values than FKM.gkb.noise.
In terms of outlier detection, the True Positive Rate (TPR) for the outlying cells shows a near perfect score for cellFCLUST and FKM.gkb.noise in both scenarios and across contamination levels, whereas F-TCLUST and FKM.noise start to stray at $5\%$, and UFTCP at $10\%$, as reported in Table 1. The same pattern occurs for the False Negative Rate (FNR), which measures the proportion of outlying cells not flagged, for cellFCLUST and FKM.gkb.noise, consistent with their better performance in clustering and parameter recovery. However, cellFCLUST is the only model capable of achieving a near-zero False Positive Rate (FPR). It is worth noting that, for the casewise models, all cells of flagged rows are considered outlying, leading to a higher number of false positives. Consequently, the reliable information discarded with the casewise models results in higher errors – also in terms of clustering and parameter estimation – even when outlying values are detected. This effect is noticeable when comparing the overall performance of cellFCLUST and FKM.gkb.noise.
Table 1: Memberships and outlier detection: proportion of Weak Assignments (WA), True and False Positive Rate (TPR and FPR), and False Negative Rate (FNR) averaged over $100$ samples per scenario, percentage of contamination, and model.
Scenario 1 Scenario 2
$\%$ out Method WA TPR FPR FNR WA TPR FPR FNR
0 cellFCLUST 0.03 - - - 0.05 - - -
F-TCLUST 0.07 - - - 0.12 - - -
FKM 0.10 - - - 0.04 - - -
FKM.noise 0.10 - - - 0.04 - - -
FKM.gkb 0.07 - - - 0.05 - - -
FKM.gkb.noise 0.09 - - - 0.08 - - -
UFTCP 0.10 - - - 0.04 - - -
1 cellFCLUST 0.03 1.00 0.00 0.00 0.05 1.00 0.00 0.00
F-TCLUST 0.00 1.00 0.25 0.00 0.01 1.00 0.24 0.00
FKM 0.18 - - - 0.13 - - -
FKM.noise 0.11 0.92 0.06 0.08 0.06 0.79 0.07 0.21
FKM.gkb 0.20 - - - 0.22 - - -
FKM.gkb.noise 0.57 1.00 0.08 0.00 0.47 1.00 0.09 0.00
UFTCP 0.17 0.93 0.09 0.07 0.14 1.00 0.09 0.00
5 cellFCLUST 0.04 0.99 0.00 0.01 0.06 0.99 0.00 0.01
F-TCLUST 1.00 0.69 0.23 0.31 1.00 0.67 0.23 0.33
FKM 0.39 - - - 0.43 - - -
FKM.noise 0.28 0.69 0.23 0.31 0.26 0.60 0.20 0.40
FKM.gkb 0.35 - - - 0.47 - - -
FKM.gkb.noise 0.22 1.00 0.36 0.00 0.17 1.00 0.37 0.00
UFTCP 0.45 0.99 0.37 0.01 0.45 1.00 0.40 0.00
10 cellFCLUST 0.05 0.99 0.00 0.01 0.07 0.99 0.00 0.01
F-TCLUST 1.00 0.47 0.23 0.53 1.00 0.46 0.23 0.54
FKM 0.66 - - - 0.70 - - -
FKM.noise 0.53 0.60 0.31 0.40 0.56 0.50 0.24 0.50
FKM.gkb 0.33 - - - 0.54 - - -
FKM.gkb.noise 0.22 0.99 0.60 0.01 0.20 0.98 0.60 0.02
UFTCP 0.60 0.81 0.45 0.19 0.65 0.80 0.45 0.20
4 Effects of the cellFCLUST tuning parameters
The proposed methodology depends on several tuning parameters: the number of clusters ($K$), the contamination level ($\alpha$), the constant for the eigenvalue-ratio constraint ($c$), and the fuzzifier ($m$). Additionally, we can also consider the scale factor ($S$), which arises when we modify $x_{ij}$ by $x_{ij}/S$ for $i=1\ldots,n$ and $j=1,\ldots J$, as a further parameter. The effect of the scale factor was already noted in [14] and is inherent when considering general scatter matrices and the definition of the membership values in (7). However, $S$ can be viewed as an actionable tuning parameter as well.
We illustrate the role of all tuning parameters using one of the samples generated in Scenario 1 of the simulation study (Section 3) with $5\%$ of contamination. It has to be highlighted that these parameters are interrelated, and a unified approach for their setting is needed. In this section, we provide tools for helping the user in selecting these parameters, which are particularly useful when no prior information is available, as is often the case in real-data applications.
@@ 4 Effects of the cellFCLUST tuning parameters
Number of clusters $K$ and contamination level $\alpha$: the number of clusters in a data set cannot be chosen independently of the level of contamination. In this framework, we evaluate the behavior of the curves representing the objective function in (4) at convergence by running cellFCLUST with different values of $K$ and $\alpha$, when $c=14$ and $m=2$. We vary $\alpha$ over the set $\{0,0.01,0.025,0.05,0.075,0.10\}$, considering that the true contamination level is $0.05$. As shown in Figure 4, the objective functions differ significantly when $\alpha=0,0.01,0.025$, since underestimating the contamination level requires additional clusters beyond those originally generated. It is worth noting that when $K=4$ and $\alpha=0$ or $0.01$, cellFCLUST finds solutions with an empty cluster, causing the algorithm to stop – the corresponding objective function values are not reported in the figure. On the other hand, when $\alpha=0.05$, which is the so generated contamination level in the data set, the objective functions for $K=2,3,4$ are close to each other, indicating that two is a suitable choice for the number of clusters. Indeed, $K=2$, which corresponds to the true value, captures the underlying structure of the data well, and increasing this number does not lead to a meaningful improvement in the objective function [33].
Figure 4: Synthetic data: objective function curves.
@@
Figure 5: Synthetic data: difference between the knee point of $\{\Delta_{ij}\}_{i=1}^{n}$ and the $\alpha$ value. The curve shows the median differences across variables as $\alpha$ varies, with $K=2$. The shaded gray area represents the variability.
@@
Figure 6: Synthetic data: $\Delta$ plots where observations are sorted according to their $\Delta$ values. Vertical, dashed, red line corresponds to the value of $\alpha$ used for the model implementation when $K=2$.
@@
We further investigate how the cellFCLUST results change with the contamination level $\alpha$ to support the previous finding by analyzing the behavior of $\{\Delta_{ij}\}_{i=1}^{n}$ for $j=1,\ldots,J$. These are computed as defined in (6), using the parameters estimated at convergence for the selected $K$. For each variable, the ordered values $\Delta_{(i)j}$, where $\Delta_{(1)j}\leq\Delta_{(2)j}\leq\ldots\leq\Delta_{(n)j}$, can be plotted. In each plot – which we discuss in detail later – we identify the knee point, corresponding to the maximum distance between each $\Delta_{(i)j}$ and the straight line connecting the endpoints $\Delta_{(1)j}$ and $\Delta_{(n)j}$, as $\alpha$ varies. Therefore, we have $J$ knee points (one per variable) for each $\alpha$. Figure 5 shows the median difference between the $J=10$ knee points and the corresponding $\alpha$, for different values of $\alpha$. The difference approaches zero when $\alpha=0.05$, with variability (gray area, calculated using the Median Absolute Deviation) close to zero, confirming the selection of this value as the appropriate contamination level for these data. Furthermore, we can verify the behavior of $\Delta_{ij}$ for each variable, recalling that they correspond to the difference between the individual contribution to the objective function when a cell is considered reliable or unreliable. In Figure 6, we plot points $\{\big((i)/n,\Delta_{(i)j}\big)\}_{i=1}^{n}$ for each variable, where observations are ordered according to their $\Delta$ values. When $\alpha=0.025$, some observations with a small value of $\Delta$ are not considered as contaminated, resulting in a lower contamination level than the theoretical one (Figure LABEL:fig:_example_Delta_2_5). The same occurs for smaller values of $\alpha$. On the other hand, too many cells than needed are flagged as unreliable when $\alpha=0.075$, as shown in Figure LABEL:fig:_example_Delta_7_5, where the values of $\Delta$ level off before the considered $\alpha$ – the same happens for $\alpha=0.10$. The appropriate choice turns out to be $0.05$ (Figure LABEL:fig:_example_Delta_5), which corresponds to the true proportion of contaminated cells in the simulated sample. Indeed, at this level of contamination, $5\%$ of the $\Delta$ values per variable are significantly lower than the majority of the others. In general, we recommend initially adopting a conservative choice of $\alpha$, then gradually decreasing it while monitoring the behavior of the objective function curves. The additional tools provided for supporting the choice of $\alpha$, such as those in Figures 5 and 6, are particularly useful when the objective function curves are smoother.
Constant $c$ for the eigenvalue-ratio constraint: the constant for the eigen-value-ratio constraint in (3) allows for different cluster shapes. When $c=1$, the clusters become spherical, and cellFCLUST with $\alpha=0$ and $p_{1}=\ldots=p_{K}$ produces similar results to those of fuzzy $c$-means. Indeed, a small value of $c$ reduces the cellFCLUST ability to detect elongated and/or differently scattered clusters, as illustrated in [14] and [34]. However, in some applications, the user may be interested in more spherical types of clusters (see, for instance, [35]).
Fuzzifier parameter $m$, scale factor $S$ and their interaction: cellFCLUST shows different behavior depending on the fuzzifier parameter and the scaling of the data. The former affects the degree of fuzziness by letting the memberships become more similar as $m$ increases. Points that are more fuzzily assigned to the clusters are usually observations located farther from the core of the clusters in the $J$-dimensional space. On the other hand, the scale factor $S$ influences the proportion of hard assignments obtained by cellFCLUST, which usually correspond to observations within the core of the clusters carrying more weight in the estimation of their parameters. This effect was previously observed in [8] and [7]. Given the property of high contrast, we thus treat the scale factor $S$ as a key tuning parameter that allows us to manage the mentioned level of hard assignment depending on the purpose of the data analysis. Furthermore, the scale factor and the fuzzifier interplay, resulting in a change in the clustering structure estimated by cellFCLUST when both $S$ and $m$ vary. Specifically, as the scale factor decreases and the fuzzifier increases, the assignments tend to become less crisp and more fuzzy. However, a complete fuzzification is not desirable from an interpretation point of view. It is worth noting that when $m=1$, the results are scale-independent since cellFCLUST returns a completely hard assignment. Since the effect of the tuning parameters $m$ and $S$ in cellFCLUST is similar to its casewise counterpart F-TCLUST, the reader can refer to [14] for further insights and numerical examples.
Restrictions on the weights $p_{k}$: one of the advantages of cellFCLUST is its ability to accommodate different cluster sizes through distinct cluster weights, in a similar manner to F-TCLUST. We do not reproduce here examples illustrating the effect of constraining $p_{k},k=1,\ldots,K$ to be equal, which the reader can find in [14]. As already mentioned, setting $p_{1}=p_{2}=\ldots=p_{K}$ forces the clusters to have roughly similar sizes. Moreover, it has to be highlighted that when $K$ is misspecificied by the user with a value larger than necessary, some cluster weights $p_{k}$ can become very close to 0, resulting in almost “empty” clusters, i.e., clusters with small $u_{ik}$ for all observations. This occurrence can reveal the potential misspecification of $K$ to the user.
The material presented in this section exemplify the role and effect of the cellFCLUST tuning parameters. All these parameters are closely related to each other, and their choice strongly depends on the application under analysis. If prior knowledge about the data set is available – for instance, regarding the expected clustering structure or the desired degree of fuzzification – this information can guide the choice of some (or, in rare cases, all) tuning parameters. When this does not occur, we suggest to select appropriate values for the tuning parameters by taking into account the purpose of the analysis. Specifically, as a first step, a simple proposal is to choose the constant $c$ by considering the order of magnitude of the ratio between the maximum and the minimum eigenvalues computed on the entire data set. Secondly, the scale factor can be set by analyzing the proportion of HA as $S$ varies, selecting a reasonable subset of values for $S$ consistent with the data under study. Indeed, although $S$ and $m$ are interconnected, the former mostly affects HA, while the latter the degree of fuzzification. In this regard, $m$ can be chosen by studying the proportion of WA. The user may prefer to select a value of $m$ that results in a proportion of WA within a specific range, depending on the goal of the analysis and the cluster configuration. Finally, the objective function curves presented in this section are a useful tool for selecting $K$ and $\alpha$, as previously described. The contamination level $\alpha$ can be deepened using the $\Delta$ plots. It is worth highlighting that the choice of $S$ and $m$ is connected to the selection of both the number of clusters and the level of contamination.